# A kernel conditional two-sample test

### v1

@@ Abstract
We propose a framework for hypothesis testing on conditional probability distributions, which we then use to construct conditional two-sample statistical tests. These tests identify the inputs — called covariates in this context — where two conditional expectations differ with high probability. Our key idea is to transform confidence bounds of a learning method into a conditional two-sample test, and we instantiate this principle for kernel ridge regression (KRR) and conditional kernel mean embeddings. We generalize existing pointwise-in-time or time-uniform confidence bounds for KRR to previously-inaccessible yet essential cases such as infinite-dimensional outputs with non-trace-class kernels. These bounds enable circumventing the need for independent data in our statistical tests, since they allow online sampling. We also introduce bootstrapping schemes leveraging the parametric form of testing thresholds identified in theory to avoid tuning inaccessible parameters, making our method readily applicable in practice. Such conditional two-sample tests are especially relevant in applications where data arrive sequentially or non-independently, or when output distributions vary with operational parameters. We demonstrate their utility through examples in process monitoring and comparison of dynamical systems. Overall, our results establish a comprehensive foundation for conditional two-sample testing, from theoretical guarantees to practical implementation, and advance the state-of-the-art on the concentration of vector-valued least squares estimation.
@@ 1 Introduction and related work
Deciding whether the inputs and outputs of two phenomena obey the same conditional relationship is essential in many areas of science. In control and robotics, one may want to detect changes in the dynamics over time or detect unusual perturbations; in industrial monitoring, to identify the performance of an equipment in varying operating conditions; and in medical studies, to compare treatment responses based on patient characteristics. These examples exhibit the common characteristic that they concern the conditional distribution of an output (next state, performance, treatment response, …) given an input parameter, which we call the covariate (current state, patient characteristics, …). While there are methods to detect global changes or mismatches between probability distributions, they generally have at least one of the following limitations: (i) they already detect a mismatch if the marginal distributions of the covariates differ while the conditional ones coincide [1]; (ii) they already detect a mismatch if they detect a global difference between the conditional distributions, disallowing questions such as at what covariate they differ or whether they differ at specific points of interest [2, 3, 4, 5, 6, 7, 8]; or (iii) they only provide guarantees in the infinitely-many samples limit [2, 3, 4, 5, 6, 7, 8, 9]. In contrast, we propose a framework enabling finite-sample guarantees for conditional hypothesis testing at arbitrary covariates, generalizing such guarantees for classical hypothesis testing in the sense of Neyman-Pearson [10, Sections 8.1–8.3.1]. To our knowledge, we are the first to formalize guarantees on the accuracy of the covariate rejection region — the set of points where the null hypothesis is rejected. This region and guarantees thereon enable identifying pointwise discrepancies between conditional distributions, rather than global discrepancies between joint or conditional distributions.
@@
Regardless of the covariate rejection region, conditional testing is notoriously hard — there exists no nontrivial test with prescribed level against all alternatives [3, 4]. This should be highly unsurprising to readers familiar with no-free-lunch theorems in statistical learning theory [11], as conditional testing typically requires making (high-probability) statements about conditional distributions at covariates where no samples were observed, and thus require extrapolation power. We provide a systematic way to escape this negative result by making explicit the ties between conditional testing and learning theory: confidence bounds of learning algorithms yield guarantees for conditional tests. This connection enables turning source conditions in statistical learning into “prior sets”, that is, classes of alternatives for which a conditional test has guarantees. To our knowledge, we are the first to formalize this connection.
@@
The natural assumption on the data in conditional testing is that outputs are sampled according to conditional distributions given inputs. In particular, outputs are not necessarily identically distributed. They may even not be independent; e.g., in the online setting [12], where data is generated sequentially as future sampling locations depend on previous measurements. This is a technical challenge for finite-sample guarantees and contrasts starkly with classical hypothesis testing [10, Chapter 8], where data is independent and identically distributed (i.i.d.) [1], approximately i.i.d. (mixing, …) [13], or at least identically distributed given previous samples (online setting) [14, 15]. A common approach in learning theory is to rely on a measure along which the covariates concentrate [16, 17, 18]. We must alleviate such assumptions, however, as we seek answers that are pointwise rather than measure-dependent. We circumvent the need for such assumptions by relying on confidence bounds of learning algorithms, that is, pointwise probabilistic bounds on the error of the learned functions. This allows us not only to provide finite-sample guarantees, but also to operate on non-i.i.d. data, whereas previous studies are confined to asymptotic tests with i.i.d. data [2, 3, 4, 5, 6, 7, 8, 9]. From a practical point of view, sequential data generation calls for repeated testing as more data becomes available, often requiring time-uniform or anytime confidence bounds [12, 19, 20, 21]. The specific test we propose is based on such a bound, which we generalize to previously inaccessible cases thanks to a new assumption we introduce and a strengthening of subgaussianity [22].
[…]
@@
Our guarantees depend on quantities common in learning theory but inaccessible in practice, such as those specifying the prior set. Furthermore, learning-theoretic bounds are commonly conservative as they are essentially worst-case in the prior set, resulting in overly cautious tests. Estimation schemes for sharper bounds are thus essential. We rely on a heuristic bootstrapping scheme we introduce to that end, leveraging the parametric structure of the testing threshold indicated by the theory.
@@
Finally, our approach is connected to multiple testing [29], sequential testing [30], anomaly detection, [31], and drift detection [32, 33]. In contrast with multiple testing, the multiple hypotheses we are simultaneously testing are potentially connected (through the prior set). Our test enables sequential testing through time-uniform bounds. Furthermore, drift detection is generally performed on marginal distributions; e.g., by testing for independence between a time process and the process of interest [32]. Finally, we illustrate how to perform anomaly detection with our tools on an example. Further exploration of these connections is an interesting avenue for future work.
@@
We provide a general framework for conditional testing and its finite-sample guarantees, and instantiate it for kernel ridge regression (KRR) together with bootstrapping schemes. Specifically:
@@
    We formalize the covariate rejection region and guarantees thereon, laying a basis for finite-sample tests (Section 3).
@@
    We turn confidence bounds of learning algorithms into conditional two-sample tests (Theorem 4.2). Our instantiation for KRR (Theorem 4.4) is the first to enjoy finite-sample guarantees, and also to allow non-i.i.d. data.
@@
  4. 4.
    We provide and implement a heuristic bootstrapping scheme of the resulting test by leveraging the parametric form of the thresholds identified in theory.
As a supplement to the main text, we provide a tutorial on using in practice the tools we develop in Appendix A, as their generality may overshadow applications. It includes summaries of the algorithms (Algorithms 1, 2, and 3) and an example on how to compare (moments of or full) conditional distributions or perform system monitoring, with numerical results. Appendix B details and evaluates numerically our bootstrapping scheme. Appendix C further elaborates on how Theorem 4.3 relates to existing bounds in the literature. This theorem is based on a more general result we show for vector-valued least squares in separable Hilbert spaces presented as a standalone result in Appendix D (specifically, Theorem D.23). Finally, Appendix E contains the proof of Theorem 4.2, and Appendix F contains further information on the numerical experiments of Section 5.
@@ 2 Preliminaries and notation
The set of positive integers is denoted by ${\mathbb{N}_{>0}}$, and for all $n\in{\mathbb{N}_{>0}}\cup\{\infty\}$, $\left[n\right]$ is the interval of integers between $1$ and $n$, with $\left[\infty\right]:={\mathbb{N}_{>0}}$. If $u$ is a sequence and $m,n\in{\mathbb{N}_{>0}}$, with $m\leq n$, we write $u_{m:n}$ for the tuple with $n-m+1$ elements $(u_{m},\dots,u_{n})$. If $m=1$, we simply write $u_{:n}$. If $\mathcal{H},\mathcal{G}$ are Hilbert, $\mathcal{L}(\mathcal{H},\mathcal{G})$ (resp., $\mathcal{L}_{\mathrm{b}}(\mathcal{H},\mathcal{G})$) is the space of linear operators (resp., bounded linear operators) from $\mathcal{H}$ to $\mathcal{G}$. If $\mathcal{G}=\mathcal{H}$, we omit it in the notations.
@@
We consider a complete probability space $(\Omega,\mathcal{A},\mathbb{P})$ and measurable spaces $(\mathcal{X},\mathcal{A}_{\mathcal{X}})$ and input and outuput measurable spaces $(\mathcal{Y},\mathcal{A}_{\mathcal{Y}})$. We assume that $\mathcal{Y}\subset\mathcal{G}$ as metric and measurable spaces, where $\mathcal{G}$ is a separable Hilbert space equipped with its Borel $\sigma$-algebra, and that $\mathcal{Y}$ is closed. We define ${\mathcal{M}_{1}^{+}}(\mathcal{Z})$ as the set of probability measures on a measurable set $\mathcal{Z}$. We identify measures on $\mathcal{Y}$ with their unique extension to $\mathcal{G}$ with support contained in $\mathcal{Y}$. Then, a Markov kernel from $\mathcal{X}$ to $\mathcal{Y}$ is a map $p:\mathcal{A}_{\mathcal{Y}}\times\mathcal{X}\to[0,1]$ such that $p(\cdot,x)$ is a probability measure on $\mathcal{Y}$ for all $x\in\mathcal{X}$ and $p(A,\cdot)$ is a measurable function for all $A\in\mathcal{A}_{\mathcal{Y}}$ [34]. If $P\in{\mathcal{M}_{1}^{+}}(\mathcal{G})$ is a probability measure such that the identity on $\mathcal{G}$ is (Bochner-)integrable w.r.t. $P$, we introduce its first moment $\mathbb{E}(P)=\int_{\mathcal{Y}}y\mathrm{d}P(y)\in\mathcal{G}$. We say that a Markov kernel $p$ from $\mathcal{X}$ to $\mathcal{Y}$ has a first moment if $p(\cdot,x)$ does for all $x\in\mathcal{X}$. Then, we use the shorthand $\mathbb{E}(p)(x):=\mathbb{E}(p(\cdot,x))$. Finally, we define the prior set $\Theta$ as a set of Markov kernels from $\mathcal{X}$ to $\mathcal{Y}$. Section 3 allows any choice of $\Theta$, and we specify in Section 4 assumptions on it under which we can provide guarantees.
@@
almost surely (a.s.), we say that $D_{p}$ is a process of transition pairs, and that $D$ is family of processes of transition pairs if (1) holds for all $p\in\Theta$. Finally, we say that $D_{p}$ is a process of independent transition pairs if (1) holds and $(X_{p,n},Y_{p,n})$ and $(X_{p,m},Y_{p,m})$ are independent for all $m,n\in{\mathbb{N}_{>0}}$ with $m\neq n$, and similarly for $D$ itself. For all $n\in{\mathbb{N}_{>0}}$, we introduce $\mathcal{D}_{n}=(\mathcal{X}\times\mathcal{Y})^{n}$ and $\mathcal{D}=\bigcup_{n=1}^{\infty}\mathcal{D}_{n}$. Elements of $\mathcal{D}$ are called data sets, and for $D:=[(x_{n},y_{n})]_{n\in\left[N\right]}\in\mathcal{D}$ we define $\mathrm{len}(D)=N$. Finally, a core assumption of our results is subgaussian noise. We consider the strengthening introduced in Mollenhauer and Schillings [22] without requiring that the variance proxy is trace-class, in contrast with the reference, as our results only require a weakening of trace-class.
@@
[…]
@@
It is well known that, for every Hermitian, positive semi-definite function $K:\mathcal{X}\times\mathcal{X}\to\mathcal{L}_{\mathrm{b}}(\mathcal{G})$, there exists a unique $\mathcal{G}$-valued RKHS of which $K$ is the unique reproducing kernel. The kernel thus fully characterizes the RKHS, and we write $(\mathcal{H},\langle\cdot,\cdot\rangle_{\mathcal{H}})=(\mathcal{H}_{K},\langle\cdot,\cdot\rangle_{K})$. We say that a kernel is trace-class if $K(x,x^{\prime})$ is, for all $(x,x^{\prime})\in\mathcal{X}$. We assume throughout that kernels are strongly measurable. We also introduce the following notion, which is new to the best of our knowledge and generalizes the common choice $K=k\cdot\mathrm{id}_{\mathcal{G}}$.
@@
We say that the kernel $K$ of a $\mathcal{G}$-valued RKHS $\mathcal{H}_{K}$ is uniform-block-diagonal (UBD) (with isometry $\iota_{\mathcal{G}}$) if there exists separable Hilbert spaces $\tilde{\mathcal{G}}$ and $\mathcal{V}$, an isometry $\iota_{\mathcal{G}}\in\mathcal{L}_{\mathrm{b}}(\tilde{\mathcal{G}}\otimes\mathcal{V},\mathcal{G})$, and a kernel ${K_{0}}:\mathcal{X}\times\mathcal{X}\to\mathcal{L}_{\mathrm{b}}(\tilde{\mathcal{G}})$ (called an elementary block) such that
@@
We often make below the assumption that ${K_{0}}$ is trace-class; this is weaker than $K$ itself being trace-class. Informally, uniform-block-diagonality is equivalent to the existence of a base of $\mathcal{G}$ where, for all $(x,x^{\prime})\in\mathcal{X}^{2}$, the operator $K(x,x^{\prime})$ has the block-diagonal, operator-valued matrix representation

### v2

# https://openreview.net/forum?id=hJKDwf32Xu

@@ Abstract
We propose a framework for hypothesis testing on conditional probability distributions, which we then use to construct statistical tests of functionals of conditional distributions. These tests identify the inputs where the functionals differ with high probability, and include tests of conditional moments or two-sample tests. Our key idea is to transform confidence bounds of a learning method into a test of conditional expectations. We instantiate this principle for kernel ridge regression (KRR) with subgaussian noise. An intermediate data embedding then enables more general tests — including conditional two-sample tests — via kernel mean embeddings of distributions. To have guarantees in this setting, we generalize existing pointwise-in-time or time-uniform confidence bounds for KRR to previously-inaccessible yet essential cases such as infinite-dimensional outputs with non-trace-class kernels. These bounds also circumvent the need for independent data, allowing for instance online sampling. To make our tests readily applicable in practice, we introduce bootstrapping schemes leveraging the parametric form of testing thresholds identified in theory to avoid tuning inaccessible parameters. We illustrate the tests on examples, including one in process monitoring and comparison of dynamical systems. Overall, our results establish a comprehensive foundation for conditional testing on functionals, from theoretical guarantees to an algorithmic implementation, and advance the state of the art on confidence bounds for vector-valued least squares estimation.
@@ 1 Introduction and related work
Deciding whether the inputs and outputs of two phenomena obey the same conditional relationship is essential in many areas of science. In control and robotics, one may want to detect changes in the dynamics over time or detect unusual perturbations; in industrial monitoring, to identify the performance of an equipment in varying operating conditions; and in medical studies, to compare treatment responses based on patient characteristics. These examples exhibit the common characteristic that they concern the conditional distribution of an output (next state, performance, treatment response, …) given an input parameter, which we call the covariate (current state, patient characteristics, …). The functional of interest is the aspect of the output distribution we care about (expectation, other moment, full distribution, …). While there are methods to detect global changes or mismatches between (functionals of) probability distributions, they generally have at least one of the following limitations: (i) they already detect a mismatch if the marginal distributions of the covariates differ while the conditional ones coincide [1]; (ii) they already detect a mismatch if they detect a global difference between the conditional distributions, disallowing questions such as at what covariate they differ or whether they differ at specific points of interest [2, 3, 4, 5, 6, 7, 8]; or (iii) they only provide guarantees in the infinitely-many samples limit [2, 3, 4, 5, 6, 7, 8, 9]. In contrast, we propose a framework enabling finite-sample guarantees for conditional hypothesis testing at arbitrary covariates, generalizing such guarantees for classical hypothesis testing in the sense of Neyman-Pearson [10, Sections 8.1–8.3.1]. To our knowledge, we are the first to formalize guarantees on the accuracy of the covariate rejection region, which is the set of points where the null hypothesis is rejected. This region and guarantees thereon enable identifying pointwise discrepancies between functionals of conditional distributions, rather than global discrepancies between joint or conditional distributions.
@@
Regardless of the covariate rejection region, conditional testing is notoriously hard — there exists no nontrivial test with prescribed level against all alternatives [3, 4]. This should be highly unsurprising to readers familiar with no-free-lunch theorems in statistical learning theory [11], as conditional testing typically requires making (high-probability) statements about conditional distributions at covariates where no samples were observed, and thus requires extrapolation power. We provide a systematic way to escape this negative result by making explicit the ties between conditional testing and learning theory: learning precedes testing, as confidence bounds of learning algorithms yield guarantees for conditional tests. This connection enables turning source conditions in statistical learning into “prior sets”, that is, classes of alternatives for which a conditional test has guarantees. To our knowledge, we are the first to formalize this connection.
@@
The natural assumption on the data in conditional testing is that outputs are sampled according to conditional distributions given inputs. In particular, outputs are not necessarily identically distributed. They may even not be independent; e.g., in the online setting [12], where data is generated sequentially as future sampling locations depend on previous measurements. This is a technical challenge for finite-sample guarantees and contrasts starkly with classical hypothesis testing [10, Chapter 8], where data is independent and identically distributed (i.i.d.) [1], approximately i.i.d. (mixing, …) [13], or at least identically distributed given previous samples (online setting) [14, 15]. A common approach in learning theory is to rely on a measure along which the covariates concentrate [16, 17, 18]. We must alleviate such assumptions, however, as we seek answers that are pointwise rather than measure-dependent. We achieve this by relying on confidence bounds of learning algorithms, that is, pointwise probabilistic bounds on the error of the learned functions. This allows us not only to provide finite-sample guarantees, but also to operate on non-i.i.d. data, whereas previous studies are confined to asymptotic tests with i.i.d. data [2, 3, 4, 5, 6, 7, 8, 9]. From a practical point of view, sequential data generation calls for repeated testing as more data becomes available, often requiring time-uniform or anytime confidence bounds [12, 19, 20, 21]. The specific test we propose is based on such a bound, which we generalize to previously inaccessible cases thanks to a new assumption we introduce and a strengthening of subgaussianity [22]. In particular, our bound is directly time-uniform.
The learning problems involved in conditional testing require mapping covariates to the functional of interest. In some cases, this may be the full conditional distribution; e.g., in two-sample or independence testing [4]. Distributional learning usually addresses this by embedding said distributions as expectations in Hilbert spaces [23, 24]. We leverage this idea by casting general conditional functional test as a comparison of conditional expectations with correspondingly-embedded data. It requires, however, confidence bounds applicable to such outputs. For instance, kernel ridge regression (KRR) admits a well-known time-uniform confidence bound for scalar outputs [12, 21], but its only extension to infinite-dimensional outputs [20] requires strong assumptions on the kernel which exclude the data embeddings that allow testing for complex functionals such as those we identify for two-sample tests. We address this by establishing a time-uniform bound that holds under significantly weaker assumptions — this is our main technical contribution. It is the first bound to hold under such general assumptions to our knowledge, and recovers exactly the ones of Abbasi-Yadkori [12] and Chowdhury and Gopalan [20] in their settings. As a result, it also enjoys the interpretation of being a frequentist correction of the Bayesian uncertainty bound obtained when interpreting KRR as a Gaussian process (GP) posterior mean [25, 26].
@@
Our guarantees depend on quantities common in learning theory but inaccessible in practice. Furthermore, learning-theoretic bounds are commonly conservative as they are worst-case in the prior set, resulting in overly cautious tests. Estimation schemes for sharper bounds are thus essential. To that end, we provide heuristic bootstrapping schemes using the parametric structure of the testing threshold identified in theory. One is adapted from Singh and Vijaykumar [27] by modifying their standard error term in accordance with our results. In principle, they can be replaced by any alternative.
@@
Finally, our approach is connected to several other methods, like multiple testing [28], though the hypotheses we test simultaneously are connected (through the prior set). Due to the uniform bounds, our test also enables sequential testing [29] and drift detection [30, 31], though the latter is generally performed on marginal distributions; e.g., by testing for independence between a time process and the process of interest. In addition, our tools can be used to perform anomaly detection [32], which we illustrate on an example, cf. Section 5. Finally, our general framework can be instantiated also to conditional independence testing, which in turn is known to be equivalent to conditional two-sample testing in some instances [4]. Further such investigations are left for future work.
@@
We provide a general framework for conditional testing and its finite-sample guarantees, and instantiate it for KRR together with bootstrapping schemes. Specifically:
@@
    We formalize the covariate rejection region and guarantees thereon, laying a basis for finite-sample conditional tests (Section 3).
@@
    We turn confidence bounds of learning algorithms into statistical tests comparing conditional expectations (Theorem 4.2). Our instantiation for KRR (Theorem 4.4) is the first to enjoy finite-sample guarantees, and also to allow non-i.i.d. data.
@@
    We specialize the KRR instantiation to a test for functionals of conditional distributions which allows conditional two-sample tests (Section 4.3).
  4. 4.
    We provide and implement heuristic bootstrapping schemes of the resulting test.
  5. 5.
@@
We expand on Contribution 3 above in Appendix A with its theoretical grounding and supporting experiments. The bootstrapping schemes and their numerical investigation are presented in Appendix B. Explicit algorithms for the test, and computing the effective test statistic and bootstrapped thresholds are in Appendix C. Appendix D further elaborates on how Theorem 4.3 relates to existing bounds in the literature. It relies on the more general Theorem E.23 on vector-valued least squares in separable Hilbert spaces, which we show and present as a standalone result in Appendix E. Appendix F is the proof of Theorem 4.2, and Appendix G has further information on our numerical experiments.
@@ 2 Preliminaries and notation
The set of positive integers is denoted by $\mathbb{N}_{>0}$}, and for all $\in{\mathbb{N}_{>0}}\cup\{\infty\}, [n]\left[n\right$] is the interval of integers between 11 and $n$, with $\left[\infty\right]:={\mathbb{N}_{>0}}$. If $u$ is a sequence and $m,n\in{\mathbb{N}_{>0}}$, with $m\leq n$, we write $u_{m:n}$ for the tuple with nm+1n-m+1 elements $(u_{m},\dots,u_{n})$. If $m=1$, we simply write $u_{:n}$. If $\mathcal{H},\mathcal{G}$ are Hilbert spaces, $\mathcal{L}(\mathcal{H},\mathcal{G})$ (resp., $\mathcal{L}_{\mathrm{b}}(\mathcal{H},\mathcal{G})$) is the space of linear operators (resp., bounded linear operators) from $\mathcal{H}$ to $\mathcal{G}$. If $\mathcal{G}=\mathcal{H}$, we simply write $\mathcal{L}(\mathcal{H})$ and $\mathcal{L}_{\mathrm{b}}(\mathcal{H})$. Finally, for a self-adjoint $A\in\mathcal{L}_{\mathrm{b}}(\mathcal{H})$, we define $\langle g,h\rangle_{A}=\langle Ag,h\rangle$, and denote the induced seminorm by $\|\cdot\|_{A}$.
@@
We consider a complete probability space $(\Omega,\mathcal{A},\mathbb{P})$ and input and output measurable spaces $(\mathcal{X},\mathcal{A}_{\mathcal{X}})$ and $(\mathcal{Y},\mathcal{A}_{\mathcal{Y}})$, respectively. We assume that $\mathcal{Y}\subset\mathcal{G}$ as metric and measurable spaces, where $\mathcal{G}$ is a separable Hilbert space equipped with its Borel $\sigma$-algebra, and that $\mathcal{Y}$ is closed. We define ${\mathcal{M}_{1}^{+}}(\mathcal{Z})$ as the set of probability measures on a measurable space $\mathcal{Z}$. We identify measures on $\mathcal{Y}$ with their unique extension to $\mathcal{G}$ with support contained in $\mathcal{Y}$. Then, a Markov kernel from $\mathcal{X}$ to $\mathcal{Y}$ is a map $p:\mathcal{A}_{\mathcal{Y}}\times\mathcal{X}\to[0,1]$ such that $p(\cdot,x)$ is a probability measure on $\mathcal{Y}$ for all $x\in\mathcal{X}$ and $p(A,\cdot)$ is a measurable function for all $A\in\mathcal{A}_{\mathcal{Y}}$ [33]. If $P\in{\mathcal{M}_{1}^{+}}(\mathcal{G})$ is a probability measure such that the identity on $\mathcal{G}$ is (Bochner-)integrable w.r.t. $P$, we introduce its first moment $\mathbb{E}(P)=\int_{\mathcal{Y}}y\mathrm{d}P(y)\in\mathcal{G}$. We say that a Markov kernel $p$ from $\mathcal{X}$ to $\mathcal{Y}$ has a first moment if $p(\cdot,x)$ does so for all $x\in\mathcal{X}$. Then, we use the shorthand $\mathbb{E}(p)(x):=\mathbb{E}(p(\cdot,x))$. Finally, we define the prior set $\Theta$ as a set of Markov kernels from $\mathcal{X}$ to $\mathcal{Y}$. Section 3 allows any such choice of $\Theta$, and we specify in Section 4 assumptions on it under which we can provide guarantees.
@@
almost surely (a.s.), we say that $D_{p}$ is a process of transition pairs, and that $D$ is family of processes of transition pairs if (1) holds for all $p\in\Theta$. Finally, we say that $D_{p}$ is a process of independent transition pairs if (1) holds and $(X_{p,n},Y_{p,n})_{n\in{\mathbb{N}_{>0}}}$ is an independent process, and similarly for $D$ itself. For all $n\in{\mathbb{N}_{>0}}$, we introduce $\mathcal{D}_{n}=(\mathcal{X}\times\mathcal{Y})^{n}$ and $\mathcal{D}=\bigcup_{n=1}^{\infty}\mathcal{D}_{n}$. Elements of $\mathcal{D}$ are called data sets, and for $D:=[(x_{n},y_{n})]_{n\in\left[N\right]}\in\mathcal{D}$ we define $\mathrm{len}(D)=N$. Finally, a core assumption of our results is subgaussian noise. We consider the strengthening introduced in Mollenhauer and Schillings [22]. In contrast to the reference, however, we do not require that the variance proxy is trace-class, as our results only require a weakening of that condition.
@@
$\mathbb{E}\left[\exp\left(\langle g,\eta_{n}\rangle_{\mathcal{G}}\right)\mid\mathcal{F}_{n}\right]\leq\exp\left(\lVert g\rVert_{R}^{2}\right),\quad\forall g\in\mathcal{G},\quad\forall n\in{\mathbb{N}_{>0}},\quad\text{\acs{as}}$ (2)
@@
It is well known that, for every Hermitian, positive semi-definite function $K:\mathcal{X}\times\mathcal{X}\to\mathcal{L}_{\mathrm{b}}(\mathcal{G})$, there exists a unique $\mathcal{G}$-valued RKHS of which $K$ is the unique reproducing kernel. The kernel thus fully characterizes the RKHS, and we write $(\mathcal{H},\langle\cdot,\cdot\rangle_{\mathcal{H}})=:(\mathcal{H}_{K},\langle\cdot,\cdot\rangle_{K})$. We say that a kernel is trace-class if $K(x,x^{\prime})$ is, for all $(x,x^{\prime})\in\mathcal{X}$. We assume throughout that kernels are strongly measurable. We also introduce the following notion, which is new to the best of our knowledge and generalizes the common choice of a diagonal kernel, $K=k\cdot\mathrm{id}_{\mathcal{G}}$.
@@
[…]
@@
In the following, we often assume that $K_{0}$} is trace-class; this is weaker than $K$ itself being trace-class. Informally, uniform-block-diagonality is equivalent to the existence of a base of $\mathcal{G}$ where, for all $(x,x^{\prime})\in\mathcal{X}^{2}$, the operator $K(x,x^{\prime})$ has the block-diagonal, operator-valued matrix representation